Determine the number of integer solutions of x1 + x2 + x3 + x4-32, where a) xi 2 0, 1 3is4 b) x1, x2 2 2, x3, X4 2 1
Answer #1

a)

x1 +x2 + x3 + x4 = 32

x1 +x2 + ..xr = n

number of integral solution =    (n+ r -1)C(r-1)

here n = 32 , r = 4

hence

(32 + 4 - 1)C(4-1)

= 35C3

= 35*34*33/6

= 6545

b)

x1 = 2 +y1 . x2 = 2 +y2 ,x3 = 1 + y3 , x4 = 1 +y4

hence

x1 +x2 + x3 + x4 = 32

2+y1 + 2 +y2 + 1 +y3 + 1 +y4 = 32 . yi >= 0

y1 +y2 +y3 +y4 = 26

= (29C3)

= (29*28*27/6)

= 3654